OpenStudy (anonymous):

A rectangular prism has length , width , and volume . What is the height of the prism? 14 m width 6m and volume 840m3

OpenStudy (anonymous):

\[V = lwh\]\[840 = 6 * 14 * h\] So solve for h: \[840 = 6 * 14 * h\]840 = 84h /84 10 = h So your height is 10:)

OpenStudy (anonymous):

10m